OpenStudy (anonymous):

Mandy obtains a $155,000 20/6 balloon mortgage with a rate of 4.25%. What will her monthly payments be? $959.81 $1044.46 $1110.47 $1225.74

OpenStudy (anonymous):

I know I use this B=B[((1+i)^(n*t)-(1+i)^(m))/((1+i)^(n*t)-(1))

OpenStudy (anonymous):

B=1550 r=0.0425 n=12 i=r/n=0.0425/12=0.0035416667 t=20 m= ?

OpenStudy (anonymous):

@ankit042

OpenStudy (ankit042):

Use the annuity(PMT) formula you identified the variables correctly

OpenStudy (ankit042):

I am getting A as the answer

OpenStudy (anonymous):

whats the PMT ? and thank you

OpenStudy (ankit042):

PMT is term used for periodic payments. If you use excel for your calculations there is a formula where you just have to put the values

OpenStudy (anonymous):

can I get the formula ?

OpenStudy (ankit042):

@NaCl you can try this http://spreadsheets.about.com/od/excelfunctions/qt/20071020.htm